OpenStudy (anonymous):

-2x + 2y = 2 -4y - 2x = -22 elimination method

Parth (parthkohli):

Subtract both equations.

OpenStudy (anonymous):

Wait Parth I think there's a correct formula on this! can u avail it here

OpenStudy (anonymous):

-2x + 2y = 2 -4y - 2x = -22

Parth (parthkohli):

What kind of formula? I think that subtracting both equations and then solving for \(y\) is enough to get \(x\).

OpenStudy (anonymous):

elimination

Parth (parthkohli):

Yep - that's elimination.